Find Delta By and Delta Cx using virtual work method. Bar areas ; Top Chords = 4; Bottom Chords = 2.5; BD = 3.

Problem 2: Find ABy and ACx. Bar areas: top chords = 4; bottom chords = 2.5; BD = 3. 40 kB 10ft 40 k 10 f T 20ft 20ft —
Answer #1

цок р .28.28! 4 Q Le 2212 O=Tau id=tant to yo) d = 26-56° Be Tum ( 0) – 26560 P = Tun() =45° EMAO ř 20lt zo Coq Sne)(40) +Virtual work method for by) Apply an unit load at B in y-direction EMA=o BK Rc Coso) (10) +Resino) (40) = 120 o AK 2054 20 RSertran P | K | Lidl AE UZ PKYAE (is) -0.2018 56.56 1.256 [28.2842| 4829000 -0.0183 40 H0.333 LOX1Z3 X 220VO -0.226. lo SPKLAt joints Efx=0 0.888 KAO 10-222 KAB cos 45 + kao Cos26:56 = 0.888 Chyzo KAB foru5+ KAD $1.26:56 = 0.222 V=45 B=2656 Еве - -о